OpenStudy (anonymous):

Use the arc length formula and the given information to find r. s = 16 cm, θ = 48°; r = ?

OpenStudy (anonymous):

what does the formula say?

OpenStudy (anonymous):

arc length

OpenStudy (anonymous):

( *2 r)/360

OpenStudy (anonymous):

60/pi cm correct?

OpenStudy (anonymous):

it says that: angle/360 * 2pier = arc length

OpenStudy (anonymous):

or wait no 120/pi cm

OpenStudy (anonymous):



OpenStudy (anonymous):

But theta must be in radian

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

48/360 * 2pie*r=16cm 48/360 * 2 pie will get 0.84 0.84r=16 16/0.84=19cm which is the radius

OpenStudy (anonymous):

First convert 48 degree into radian

OpenStudy (anonymous):

4pi/15

OpenStudy (anonymous):

yes then 16/4pi/15

OpenStudy (anonymous):